# A Connection Between Covariate Adjustment and Stratified Randomization in Randomized Clinical Trials

### v2

@@ A Connection Between Covariate Adjustment and Stratified Randomization in Randomized Clinical Trials
The statistical efficiency of randomized clinical trials can be improved by incorporating information from baseline covariates (i.e., pre-treatment patient characteristics). This can be done in the design stage using stratified (permutated block) randomization or in the analysis stage through covariate adjustment. This article makes a connection between covariate adjustment and stratified randomization in a general framework where all regular, asymptotically linear estimators are identified as augmented estimators. From a geometric perspective, covariate adjustment can be viewed as an attempt to approximate the optimal augmentation function, and stratified randomization improves a given approximation by moving it closer to the optimal augmentation function. The efficiency benefit of stratified randomization is asymptotically equivalent to attaching an optimal augmentation term based on the stratification factor. Under stratified randomization, adjusting for the stratification factor only in data analysis is not expected to improve efficiency, and the key to efficient estimation is incorporating new prognostic information from other covariates. In designing a trial with stratified randomization, it is not essential to include all important covariates in the stratification, because their prognostic information can be incorporated through covariate adjustment. These observations are confirmed in a simulation study and illustrated using real clinical trial data.
@@ 1 Introduction
In this article, we make a general connection between covariate adjustment and stratified randomization and discuss its practical implications. Under the aforementioned augmentation approach, which covers virtually all estimators of practical interest, we note that covariate adjustment can be viewed geometrically as an attempt to approximate the optimal augmentation function in a suitable function space. From this geometric perspective, stratified randomization can be interpreted as a way to improve a given approximation by projecting it into a certain subspace containing the optimal augmentation function. The efficiency benefit of stratified randomization is asymptotically equivalent to adding an optimal augmentation term based on the stratification factor. This connection helps characterize efficient estimators under stratified randomization and has important practical implications:
  • Under stratified randomization, adjusting for the stratification factor only in the analysis of trial data is not expected to improve statistical efficiency. Adjusting for the stratification factor is common practice in clinical trials and consistent with regulatory guidelines 24, 25, 26, 1. However, the key to efficient estimation is incorporating new prognostic information from other baseline covariates than the stratification factor (whose prognostic information has already been utilized in stratified randomization). Such covariates may include continuous covariates underlying the stratification factor as well as additional covariates unrelated to the stratification factor.
@@
The rest of the article is organized as follows. In Section 2, we formulate the estimation problem and describe treatment effect estimators including augmented estimators. In Section 3, we summarize asymptotic results for stratified randomization, make a connection with covariate adjustment, and discuss its practical implications. A simulation study is reported in Section 4, followed by a real example in Section 5. The article concludes with a discussion in Section 6. Technical proofs are provided in Appendix.
2 Covariate Adjustment
@@ 2.1 Estimation Problem
We start by formulating the estimation problem using potential outcomes 27. For a generic patient in the target population, let $Y(a)$ denote the potential outcome for treatment $a$, where $a=1$ for an experimental treatment and $a=0$ for a control treatment, which may be placebo or standard of care. For each $a\in\{0,1\}$, we assume $\operatorname{E}\{Y(a)^{2}\}<\infty$ and denote $\mu_{a}=\operatorname{E}\{Y(a)\}$. Suppose the treatment effect of interest is $\delta=g(\mu_{1})-g(\mu_{0})$, where $g$ is a smooth and strictly increasing function specified by the investigator. Popular choices for $g$ include identity (for continuous and other types of outcomes), logarithm (for outcomes with positive means) and logit (for binary outcomes).
@@
2.2 Estimators
While stratified randomization is of particular interest in this article, simple randomization provides a natural framework for introducing estimators. Therefore, in the rest of this section, we assume that simple randomization is used to make treatment assignments. This means that the $A_{i}$’s are independent of each other and of the $\boldsymbol{Z}_{i}$’s, with $\operatorname{P}(A_{i}=1)\equiv\pi\in(0,1)$. It follows that the $\boldsymbol{O}_{i}$’s are independent copies of $\boldsymbol{O}=(\boldsymbol{W},A,Y)$, where $A$ is Bernoulli with $\operatorname{P}(A=1|\boldsymbol{Z})=\operatorname{P}(A=1)=\pi$ and $Y=Y(A)=AY(1)+(1-A)Y(0)$.
@@ 2.2 Estimators
There are many covariate-adjusted estimators of $\delta$ that incorporate covariate information for improved efficiency. Here we describe one such estimator based on a standard regression model:
$\operatorname{E}(Y|A,\boldsymbol{V})=h\left((1,A,\boldsymbol{V}^{\text{T}},A\boldsymbol{V}^{\text{T}})\mbox{\boldmath${\beta}$}\right),$ (1)
where $\boldsymbol{V}$ is a vector-valued function of $\boldsymbol{W}$, $h$ a specified inverse link function, and ${\beta}$ an unknown parameter vector. The interaction term in model (1) is optional in practice but necessary in the subsequent discussion. Suppose ${\beta}$ is estimated by solving the estimating equation
$\sum_{i=1}^{n}\left\{Y_{i}-h(1,A_{i},\boldsymbol{V}_{i}^{\text{T}},A_{i}\boldsymbol{V}_{i}^{\text{T}})\mbox{\boldmath${\beta}$}\right\}(1,A_{i},\boldsymbol{V}_{i}^{\text{T}},A_{i}\boldsymbol{V}_{i}^{\text{T}})^{\text{T}}=\boldsymbol{0}.$ (2)
[…]
Examples of the regression estimator include the analysis of covariance (ANCOVA) estimator for continuous outcomes 3 and the standardized logistic regression estimator for binary outcomes 6. It is worth noting that, whether model (1) is correct or not, $\widehat{\delta}_{\text{reg}}$ is consistent for $\delta$ and asymptotically normal with asymptotic variance $\operatorname{Var}\{\psi(\boldsymbol{O})-(A-\pi)b_{\text{reg}}(\boldsymbol{W})\}$, where
[…]
For full generality, we will consider a large class of augmented estimators of $\delta$ that encompass all estimators of practical interest. The augmentation approach is described in detail by the original authors 3, 4 and summarized by Zhang, Zhang and Liu 29 using notations similar to ours. In this paragraph, we give a high-level description of this approach, omitting technical details and nuances. An augmented estimator of $\delta$ may be obtained as $\widehat{\delta}_{\text{aug}}(b)=\widehat{\delta}_{\text{emp}}-n^{-1}\sum_{i=1}^{n}(A_{i}-\pi)b(\boldsymbol{W}_{i})$, where $b(\boldsymbol{W})$ is a real-valued function of $\boldsymbol{W}$ with finite variance. For any fixed $b$, $\widehat{\delta}_{\text{aug}}(b)$ is consistent and asymptotically normal with asymptotic variance $\sigma^{2}(b)=\operatorname{Var}\{\psi(\boldsymbol{O})-(A-\pi)b(\boldsymbol{W})\}$. Furthermore, any estimator of $\delta$ with certain desirable properties must be asymptotically equivalent to $\widehat{\delta}_{\text{aug}}(b)$ for some $b$ (and thus have the same asymptotic variance). For example, the regression estimator $\widehat{\delta}_{\text{reg}}$ is asymptotically equivalent to $\widehat{\delta}_{\text{aug}}(b_{\text{reg}})$ with the same asymptotic variance $\sigma^{2}(b_{\text{reg}})$. For efficient estimation of $\delta$, it suffices to restrict attention to augmented estimators and search for the most efficient one with the smallest asymptotic variance. The optimal choice of $b$, which minimizes $\sigma^{2}(b)$, is given by
@@
where $m_{a}(\boldsymbol{W})=\operatorname{E}(Y|\boldsymbol{W},A=a)=\operatorname{E}\{Y(a)|\boldsymbol{W}\}$, $a=0,1$. Although $b_{\text{opt}}$ is unknown, it can be estimated from trial data using a regression model for $\operatorname{E}(Y|A,\boldsymbol{W})$ or a machine learning algorithm. Let $\widehat{b}$ be a generic estimator of $b_{\text{opt}}$ and assume that $\widehat{b}$ converges to some limit function $b$, which may be different from $b_{\text{opt}}$. Under quite general conditions, $\widehat{\delta}_{\text{aug}}(\widehat{b})$ (or a cross-fitted version of it) is asymptotically equivalent to $\widehat{\delta}_{\text{aug}}(b)$ with the same asymptotic variance $\sigma^{2}(b)$. This result holds regardless of the convergence rate of $\widehat{b}$, which supports the use of nonparametric machine learning methods to estimate $b_{\text{opt}}$.
Remark 1.
Covariate adjustment can be viewed geometrically as an attempt to approximate the optimal augmentation function $b_{\text{opt}}$. Not only do we know that $b_{\text{opt}}$ minimizes the asymptotic variance $\sigma^{2}(b)$ of an augmented estimator, we can show that the value of $\sigma^{2}(b)$ is directly related to the mean square (i.e., $L_{2}$) distance between $b$ and $b_{\text{opt}}$. Write $\|d\|_{2}^{2}=\operatorname{E}\{d(\boldsymbol{W})^{2}\}$ for any real-valued function $d$ of $\boldsymbol{W}$. In Appendix, we show that, for any augmentation function $b$,
In other words, the extra variance of a sub-optimal estimator, $\sigma^{2}(b)-\sigma^{2}(b_{\text{opt}})$, is proportional to the approximation error $\|b-b_{\text{opt}}\|_{2}^{2}$. The practical implication is that, while it may be unrealistic to hope that $b=b_{\text{opt}}$, it is generally helpful to reduce the approximation error $\|b-b_{\text{opt}}\|_{2}^{2}$.
3 Stratified Randomization
3.1 Description and Known Results
@@
3.2 Connection with Covariate Adjustment
In this subsection, we give a geometric interpretation to equation (4) that helps make a connection with covariate adjustment. We will show that the variance reduction effect of stratified randomization is equivalent to moving a given augmentation function $b$ closer to $b_{\text{opt}}$ under simple randomization. Recall from Remark 1 that, under simple randomization, a shorter distance to $b_{\text{opt}}$ leads to a smaller asymptotic variance.
Some notations are necessary for describing the result. We write $\mathcal{B}$ for the space of augmentation functions, i.e., =$\mathcal{B}=\{b:\|b\|_{2}<\infty\}$. This will be regarded as a Hilbert space with inner product $\langle b_{1},b_{2}\rangle=\operatorname{E}\{b_{1}(\boldsymbol{W})b_{2}(\boldsymbol{W})\}$. Let $\mathcal{C}$ denote the linear subspace of $\mathcal{B}$ consisting of functions of $S$; this can be formally defined as $\mathcal{C}=\{c\in\mathcal{B}:\operatorname{Var}\{c(\boldsymbol{W})|S\}\equiv 0\}$. Let $\mathcal{C}^{\perp}$ denote the orthogonal complement of $\mathcal{C}$ in $\mathcal{B}$, so $\mathcal{C}^{\perp}=\{b\in\mathcal{B}:\operatorname{E}\{b(\boldsymbol{W})|S\}\equiv 0\}$. These subspaces are illustrated in Figure 1. Let the operator $\Lambda$ be defined by $\Lambda b=b_{\text{opt}}+\Pi(b-b_{\text{opt}}|\mathcal{C}^{\perp})$, where $\Pi(b-b_{\text{opt}}|\mathcal{C}^{\perp})$ is the projection of $b-b_{\text{opt}}$ into $\mathcal{C}^{\perp}$. As shown in Figure 1, $\Lambda b$ is the result of projecting $b$ into the affine subspace $b_{\text{opt}}+\mathcal{C}^{\perp}$.
Figure 1: A geometric interpretation of stratified randomization (see Section 3.2 for definitions).
In Appendix, we show that, for any $b\in\mathcal{B}$,
@@ 3.2 Connection with Covariate Adjustment
Thus, for an estimator with (asymptotic) augmentation function $b$, stratified randomization amounts to replacing $b$ with $\Lambda b$ under simple randomization. Because $\Lambda$ projects into a subspace containing $b_{\text{opt}}$, $\Lambda b$ is generally closer to $b_{\text{opt}}$ than $b$ is. By Remark 1, this implies $\sigma^{2}(\Lambda b)\leq\sigma^{2}(b)$.
Remark 2.
Equation (5) makes it clear that $\sigma_{\text{st}}^{2}(b)$ reaches its minimum, $\sigma^{2}(b_{\text{opt}})$, when and only when $b-b_{\text{opt}}\in\mathcal{C}$. In other words, $\sigma_{\text{st}}^{2}(b)$ is minimized by taking $b(\boldsymbol{W})=b_{\text{opt}}(\boldsymbol{W})+c(S)$ for an arbitrary function $c$. This result indicates that stratified randomization provides some latitude in efficient estimation by allowing $b$ to differ from $b_{\text{opt}}$ by a function of $S$. The fact that $\sigma_{\text{st}}^{2}(b)$ and $\sigma^{2}(b)$ have the same minimum value is consistent with the main result of Rafi 30, which states that, for estimating the average treatment effect, the same efficiency bound holds under CAR and simple randomization.
Remark 3.

### v3

@@ A Connection Between Covariate Adjustment and Stratified Randomization in Randomized Clinical Trials
The statistical efficiency of randomized clinical trials can be improved by incorporating information from baseline covariates (i.e., pre-treatment patient characteristics). This can be done in the design stage using stratified (permutated block) randomization or in the analysis stage through covariate adjustment. This article makes a connection between covariate adjustment and stratified randomization in a general framework where all regular, asymptotically linear estimators are identified as augmented estimators. From a geometric perspective, covariate adjustment can be viewed as an attempt to approximate the optimal augmentation function, and stratified randomization improves a given approximation by moving it closer to the optimal augmentation function. The efficiency benefit of stratified randomization is asymptotically equivalent to attaching an optimal augmentation term based on the stratification factor. In designing a trial with stratified randomization, it is not essential to include all important covariates in the stratification, because their prognostic information can be incorporated through covariate adjustment. Under stratified randomization, adjusting for the stratification factor only in data analysis is not expected to improve efficiency, and the key to efficient estimation is incorporating prognostic information from all important covariates. These observations are confirmed in a simulation study and illustrated using real clinical trial data.
@@ 1 Introduction
In this article, we make a general connection between covariate adjustment and stratified randomization and discuss its statistical and practical implications. Under the aforementioned augmentation approach, which covers virtually all estimators of practical interest, we note that covariate adjustment can be viewed geometrically as an attempt to approximate the optimal augmentation function in a suitable function space. From this geometric perspective, stratified randomization can be interpreted as a way to improve a given approximation by projecting it into a certain subspace containing the optimal augmentation function. The efficiency benefit of stratified randomization is asymptotically equivalent to adding an optimal augmentation term based on the stratification factor. This connection helps characterize efficient estimators under stratified randomization and has important practical implications:
@@
  • Under stratified randomization, adjusting for the stratification factor only in the analysis of trial data is not expected to improve statistical efficiency. Adjusting for the stratification factor is common practice in clinical trials and consistent with regulatory guidelines 28, 29, 30, 1. However, the key to efficient estimation is incorporating prognostic information from all important baseline covariates. Such covariates may include continuous covariates underlying the stratification factor as well as additional covariates unrelated to the stratification factor.
@@
The rest of the article is organized as follows. In Section 2, we formulate the estimation problem, characterize covariate-adjusted treatment effect estimators, and summarize asymptotic results for stratified randomization. In Section 3, we make a connection between covariate adjustment and stratified randomization, and discuss its statistical and practical implications. A simulation study is reported in Section 4, followed by a real example in Section 5. The article concludes with a discussion in Section 6. Appendix A describes a specific type of covariate-adjusted estimators and Appendix B provides proofs of theoretical results.
2 Preliminaries
@@ 2.1 Estimation Problem
We start by formulating the estimation problem using potential outcomes 31. For a generic patient in the target population, let $Y(a)$ denote the potential outcome for treatment $a$, where $a=1$ for an experimental treatment and $a=0$ for a control treatment, which may be placebo or standard of care. For each $a\in\{0,1\}$, we assume $\operatorname{E}\{Y(a)^{2}\}<\infty$ and denote $\mu_{a}=\operatorname{E}\{Y(a)\}$. Suppose the treatment effect of interest is $\delta=g(\mu_{1})-g(\mu_{0})$, where $g$ is a smooth and strictly increasing function specified by the investigator. Popular choices for $g$ include identity (for continuous and other types of outcomes), logarithm (for outcomes with positive means) and logit (for binary outcomes).
@@
2.2 Covariate Adjustment
While stratified randomization is of particular interest in this article, simple randomization provides a natural framework for introducing estimators. Therefore, in the rest of this subsection, we assume that simple randomization is used to make treatment assignments. This means that the $A_{i}$’s are independent of each other and of the $\boldsymbol{Z}_{i}$’s, with $\operatorname{P}(A_{i}=1)\equiv\pi\in(0,1)$. It follows that the $\boldsymbol{O}_{i}$’s are independent copies of $\boldsymbol{O}=(\boldsymbol{W},A,Y)$, where $A$ is Bernoulli with $\operatorname{P}(A=1|\boldsymbol{Z})=\operatorname{P}(A=1)=\pi$ and $Y=Y(A)=AY(1)+(1-A)Y(0)$.
@@ 2.2 Covariate Adjustment
The information in $\boldsymbol{W}$ can be incorporated using an augmentation approach 3, 4. In the present setting, an augmented estimator of $\delta$ may be obtained as $\widehat{\delta}_{\text{aug}}(b)=\widehat{\delta}_{\text{emp}}-n^{-1}\sum_{i=1}^{n}(A_{i}-\pi)b(\boldsymbol{W}_{i})$, where $b(\boldsymbol{W})$ is a real-valued function of $\boldsymbol{W}$ such that $\|b\|_{2}^{2}:=\operatorname{E}\{b(\boldsymbol{W})^{2}\}<\infty$. The set of all such functions is denoted by $\mathcal{B}$ and regarded as an $L_{2}$ space. For any fixed $b\in\mathcal{B}$, $\widehat{\delta}_{\text{aug}}(b)$ is consistent and asymptotically normal with asymptotic variance $\sigma^{2}(b)=\operatorname{Var}\{\psi(\boldsymbol{O})-(A-\pi)b(\boldsymbol{W})\}$. Furthermore, any estimator of $\delta$ that is regular and asymptotically linear must be asymptotically equivalent to $\widehat{\delta}_{\text{aug}}(b)$ for some $b$ (and thus have the same asymptotic variance) 13. The optimal choice of $b$, which minimizes $\sigma^{2}(b)$, is given by
@@
where $m_{a}(\boldsymbol{W})=\operatorname{E}(Y|\boldsymbol{W},A=a)=\operatorname{E}\{Y(a)|\boldsymbol{W}\}$, $a=0,1$.
Although $b_{\text{opt}}$ is unknown, it can be estimated from trial data using a regression model for $\operatorname{E}(Y|A,\boldsymbol{W})$ 3 or a machine learning algorithm 10. Let $\widehat{b}$ be a generic estimator of $b_{\text{opt}}$ and assume that $\widehat{b}$ converges to some limit function $b\in\mathcal{B}$, which may be different from $b_{\text{opt}}$, in the sense that $\|\widehat{b}-b\|_{2}=o_{p}(1)$. Under quite general conditions, $\widehat{\delta}_{\text{aug}}(\widehat{b})$ is asymptotically equivalent to $\widehat{\delta}_{\text{aug}}(b)$ with the same asymptotic variance $\sigma^{2}(b)$. This is the case if $\widehat{b}$ belongs to a Donsker class 24 or if $\widehat{\delta}_{\text{aug}}(\widehat{b})$ employs sample splitting 25, 26, 27. We will not be concerned about these technical issues in this article and will simply assume that $\widehat{b}$ and $\widehat{\delta}_{\text{aug}}(\widehat{b})$ possess the properties stated in this paragraph.
In Appendix A, we describe a class of regression estimators of $\delta$ that are commonly used in practice. These regression estimators are not constructed as, but are asymptotically equivalent to, augmented estimators.
2.3 Stratified Randomization
@@
3 Results and Implications
3.1 Geometric Interpretations
In this subsection, we make a connection between covariate adjustment and stratified randomization based on their geometric interpretations. The geometric arguments we use to make the connection are helpful but not essential for understanding the statistical and practical implications of the connection (to be discussed in Sections 3.2 and 3.3).
We note first that covariate adjustment can be viewed geometrically as an attempt to approximate the optimal augmentation function $b_{\text{opt}}$. Not only do we know that $b_{\text{opt}}$ minimizes the asymptotic variance $\sigma^{2}(b)$ of an augmented estimator, we can show that the value of $\sigma^{2}(b)$ is directly related to the mean square (i.e., $L_{2}$) distance between $b$ and $b_{\text{opt}}$.
Proposition 1.
For any $b\in\mathcal{B}$, we have
In words, the extra asymptotic variance of a sub-optimal estimator, $\sigma^{2}(b)-\sigma^{2}(b_{\text{opt}})$, is proportional to the approximation error $\|b-b_{\text{opt}}\|_{2}^{2}$. Proposition 1 indicates that, while it may be unrealistic to hope that $b=b_{\text{opt}}$, it is generally helpful to reduce the approximation error $\|b-b_{\text{opt}}\|_{2}^{2}$.
Next, we give a geometric interpretation to equation (1) that helps make a connection with covariate adjustment. We will show that the variance reduction effect of stratified randomization is equivalent to moving a given augmentation function $b$ closer to $b_{\text{opt}}$ under simple randomization. Some notations are necessary for describing this result. Recall that $\mathcal{B}$, the space of augmentation functions, is a Hilbert space with inner product $\langle b_{1},b_{2}\rangle=\operatorname{E}\{b_{1}(\boldsymbol{W})b_{2}(\boldsymbol{W})\}$. Figure 1 visualizes the space $\mathcal{B}$ as a plane, where each individual point represents an augmentation function. Let $\mathcal{C}$ denote the linear subspace of $\mathcal{B}$ consisting of functions of $S$; this is illustrated in Figure 1 as the horizontal line passing through 0. Let $\mathcal{C}^{\perp}$ denote the orthogonal complement of $\mathcal{C}$ in $\mathcal{B}$, consisting of elements of $\mathcal{B}$ that are orthogonal to (i.e., have zero inner product with) all elements of $\mathcal{C}$. As a basic fact in functional analysis, the difference $b-b_{\text{opt}}$ can be decomposed as
$b-b_{\text{opt}}=\Pi(b-b_{\text{opt}}|\mathcal{C})+\Pi(b-b_{\text{opt}}|\mathcal{C}^{\perp}),$
where $\Pi$ denotes projection (e.g., $\Pi(b-b_{\text{opt}}|\mathcal{C})$ is the unique point in $\mathcal{C}$ that is closest to $b-b_{\text{opt}}$). Orthogonality implies that
$\|b-b_{\text{opt}}\|_{2}^{2}=\|\Pi(b-b_{\text{opt}}|\mathcal{C})\|_{2}^{2}+\|\Pi(b-b_{\text{opt}}|\mathcal{C}^{\perp})\|_{2}^{2}.$
Let $\Lambda b=b_{\text{opt}}+\Pi(b-b_{\text{opt}}|\mathcal{C}^{\perp})$, which is closer to $b_{\text{opt}}$ than $b$ is because
$\|\Lambda b-b_{\text{opt}}\|_{2}=\|\Pi(b-b_{\text{opt}}|\mathcal{C}^{\perp})\|_{2}\leq\|b-b_{\text{opt}}\|_{2}.$
Figure 1 illustrates that $\Lambda b$ is the result of projecting $b$ into the affine subspace $b_{\text{opt}}+\mathcal{C}^{\perp}$, which is shown in Figure 1 as the vertical line passing through $b_{\text{opt}}$. In other words, the operator $\Lambda$ moves a given $b\in\mathcal{B}$ in parallel to $\mathcal{C}$ to minimize its distance from $b_{\text{opt}}$. Its relationship with stratified randomization is given in the next result.
Proposition 2.
For any $b\in\mathcal{B}$, we have
@@ 3.1 Geometric Interpretations
Thus, for an estimator with (asymptotic) augmentation function $b$, stratified randomization amounts to replacing $b$ with $\Lambda b$ under simple randomization, equating $\sigma_{\text{st}}^{2}(b)$ to $\sigma^{2}(\Lambda b)$. Because $\Lambda b$ is closer to $b_{\text{opt}}$ than $b$ is, Proposition 1 further implies than $\sigma^{2}(\Lambda b)\leq\sigma^{2}(b)$.
Figure 1: A geometric interpretation of stratified randomization. The plane represents $\mathcal{B}$, the space of augmentation functions. The horizontal line passing through 0 represents the subspace $\mathcal{C}$ consisting of functions of $S$. The vertical line passing through $b_{\text{opt}}$ represents the affine subspace $b_{\text{opt}}+\mathcal{C}^{\perp}$. The dotted horizontal line segment represents the operator $\Lambda$, which projects a given $b\in\mathcal{B}$ into the affine space $b_{\text{opt}}+\mathcal{C}^{\perp}$. The fact that $\Lambda b$ is closer to $b_{\text{opt}}$ than $b$ is explains the variance reduction effect of stratified randomization. The point $\Lambda 0=c_{\text{opt}}$ is the optimal augmentation function based on $S$. See Section 3.1 for more details.
3.2 Statistical Implications
3.2.1 Efficient estimation
Equation (3) makes it clear that $\sigma_{\text{st}}^{2}(b)$ reaches its minimum, $\sigma^{2}(b_{\text{opt}})$, when and only when $b-b_{\text{opt}}\in\mathcal{C}$. In other words, $\sigma_{\text{st}}^{2}(b)$ is minimized by taking $b(\boldsymbol{W})=b_{\text{opt}}(\boldsymbol{W})+c(S)$ for an arbitrary function $c$. This result indicates that stratified randomization provides some latitude in efficient estimation by allowing $b$ to differ from $b_{\text{opt}}$ by a function of $S$. More importantly, it confirms that the efficient estimation strategy based on $\widehat{\delta}_{\text{aug}}(\widehat{b})$, with $\widehat{b}$ targeting $b_{\text{opt}}$, remains appropriate under stratified randomization.
The fact that $\sigma_{\text{st}}^{2}(b)$ and $\sigma^{2}(b)$ have the same minimum value is consistent with the main result of Rafi 33, which states that, for estimating the average treatment effect, the same efficiency bound holds under CAR and simple randomization. Thus, at least for $\delta=\mu_{1}-\mu_{0}$, the class of augmented estimators is large enough to attain the efficiency bound under stratified randomization, even though it is not known (by the author) to contain (up to asymptotic equivalence) all regular, asymptotically linear estimators under stratified randomization.
3.2.2 Analytical analogue of stratified randomization